OpenStudy (anonymous):

Find the product 5 10 -x--= 9 11

OpenStudy (anonymous):

\[\Large \frac{5}{9} \times \frac{{10}}{{11}} = \frac{{5 \times 10}}{{9 \times 11}} = \frac{?}{?}\]

OpenStudy (anonymous):